OpenStudy (anonymous):

solve.Three times the difference of 14 subtracted from a certain number is 3 less than the number increased by 7. Find the number.

OpenStudy (anonymous):

3(x-14) = (x+7)-3 3x-42 = x+4 2x = 46 x = 23 ----> answer :-)